Suppose the price ratio of two goods is 3/4 and Shubman, another student in college has a budget of $100. If the price of good X increases from 6 to 12, what is the new price

ratio? Show your work


1
Expert's answer
2021-10-07T16:14:01-0400

Solution:

The price ratio = the slope of the budget

The budget constraint: PxX + PyY = I

Price ratio = "\frac{3}{4}"

Initial price of good X (Px) = 6

Price ratio = "\frac{Px}{Py}"

"\frac{3}{4} = \frac{6}{Py}"


"3 = \frac{24}{Py}"


Py = 8


6X + 8Y = 100

The initial price of good Y = 8


Increase in the price of good X to 12:


The new price ratio = "\frac{Px}{Py} = \frac{12}{8} = \frac{3}{2} = 1.5"


The new price ratio = "\frac{3}{2} = 1.5"